OpenStudy (anonymous):

You plan a car trip for which you want to average 90 km/h. You cover the first half of the distance at an average speed of only 48 km/h. What must your average speed be in the second half of the trip to meet your goal?

OpenStudy (anonymous):

Oh wait. replace the 48 with 58. Sorry.

OpenStudy (anonymous):

guess we have to use \(\frac{D}{T}=R\)

OpenStudy (anonymous):

notice that one thing you are not told is the distance travelled since you are not told, it obviously doesn't matter, so we have a choice, we can use the variable D to compute with or (what is almost always easier) is to pick a number, compute with that number, and whatever answer we get will be correct

OpenStudy (anonymous):

which would you prefer?

OpenStudy (anonymous):

Pick a number.

OpenStudy (anonymous):

ok i pick 100

OpenStudy (anonymous):

your time so far is \(\frac{50}{58}\) i.e. distance divide by rate

OpenStudy (anonymous):

the rest of the time will be \(\frac{50}{x}\) the rate you do not know

OpenStudy (anonymous):

and you must have the total distance divided by the total time be 90 i.e. \[\frac{100}{\frac{50}{58}+\frac{50}{x}}=90\] solve for \(x\)

OpenStudy (anonymous):

you need help with that or are you good?

OpenStudy (anonymous):

I talked it over with some friends also doing this homework and we still dont get it... FYI This is Physics Homework.

OpenStudy (anonymous):

Oh wait we got it thanks anyway!